An old jar of dimes and quarters contains 200 coins total, and the value of all the coins is $42.50. How many quarters are in the jar?
Answer by fractalier(6550)

Call the numbers of dimes and quarters, d and q. Then we have
d + q = 200 and the value equation becomes
10d + 25q = 4250 (we work in cents)
Now multiply the first equation by ten and subtract...
10d + 25q = 4250
-(10d + 10q = 2000)
------------------
15q = 2250
q = 150 quarters
d = 50 dimes